If 1/4 represents 40% of a number, what is the value of that original number?
Found 2 solutions by dabanfield, richwmiller:
Answer by dabanfield(803)

If 1/4 represents 40% of a number, what is the value of that original number?
Let x be the original number.
Then we have:
1/4 = .40*x
Solve for x.

Answer by richwmiller(17219)

40%=.40
.40x=1/4
.40x=.25
40x=25
8x=5
x=5/8=.625